OpenStudy (anonymous):

find the product. (8z + 3) (5z +4)

OpenStudy (muhammad_nauman_umair):

=8z( 5z + 4 ) +3 ( 5z + 4 )

OpenStudy (anonymous):



OpenStudy (anonymous):

@kayla.

OpenStudy (texaschic101):

FOIL method....you familiar with that ?

OpenStudy (anonymous):

who

OpenStudy (texaschic101):

FOIL stands for first,outside,inside,last (8z + 3)(5z + 4) First -- multiply first terms in each set....8z * 5z = 40z^2 Outside -- multiply outside terms in each set ... 8z * 4 = 32x Inside -- multiply inside terms in each set .. 3 * 5z = 15z last -- multiply last terms in each set ... 3 * 4 = 12 Now we have : 40z^2 + 32z + 15z + 12 -- combine like terms 40z^2 + 47z + 12 ===> your answer

OpenStudy (texaschic101):

any questions ?